# 3AM: 3egment Anything with Geometric Consistency in Videos

### v3

@@ 4.1.2 Replica
SAM2Long [15] 0.7691 0.5195 0.6273
DAM4SAM [56] 0.7744 0.5135 0.6124
3AM (Ours) 0.8119 0.6381 0.6793
On the Replica dataset, 3AM again achieves the best performance across all three evaluation metrics as shown in Tab. 4. Compared to SAM2, which attains an IoU of 0.4424, SAM2Long (0.7691) and DAM4SAM (0.7744) provide substantial improvements due to their enhanced memory-selection strategies. However, 3AM surpasses all prior methods, reaching an IoU of 0.8119.
The gains are even more pronounced on the metrics that isolate visible-object performance. For Positive IoU, 3AM obtains 0.6381, clearly higher than SAM2Long (0.5195) and DAM4SAM (0.5135). Similarly, 3AM achieves the highest Successful IoU at 0.6793, indicating stronger consistency in frames where the object is correctly localized. These results confirm 3AM provides robust object tracking under the large viewpoint variations characteristic of the Replica dataset.
These results demonstrate that 3AM consistently surpasses prior methods and is particularly effective in scenarios involving object reappearance and large viewpoint variation, despite not relying on the explicit memory-selection mechanisms used in previous approaches. Due to space limit, more visualization is provided in the supplementary materials.
@@ 4.3 Ablation Study
Memroy Selection.
Table 6: Video Object Segmentation results between different memory selection methods on ScanNet++ Whole Set. Positive IoU is computed over frames with ground-truth masks. Successful IoU is computed over frames where a ground-truth object exists and the IoU of prediction and GT $\neq 0$.
Method IoU $\uparrow$ Positive IoU $\uparrow$ Successful IoU $\uparrow$
3AM 0.8898 0.5630 0.7155
DAM4SAM-3AM [56] 0.8941 0.5471 0.7204
SAM2Long-3AM [15] 0.9004 0.5498 0.7361
We further analyze the role of memory selection by combining 3AM with existing strategies proposed for SAM2. Table 6 reports results on the ScanNet++ Whole Set. 3AM, which uses the original SAM2 memory-selection mechanism, already achieves strong performance with an IoU of 0.8898. When we incorporate alternative memory-selection policies, such as those from DAM4SAM [56] or SAM2Long [15], we observe modest improvements: DAM4SAM-3AM slightly increases Successful IoU to 0.7204, while SAM2Long-3AM achieves the highest IoU (0.9004) and Successful IoU (0.7361). These results suggest that 3AM already delivers strong and stable performance without relying on specialized memory-selection policies. While alternative strategies such as those from SAM2Long or DAM4SAM offer small additional gains, their impact is relatively limited compared to the improvements brought by 3AM itself. We view the design of memory-selection mechanisms tailored specifically for 3AM as an interesting direction for future work.
3D Foundataion Models.
@@
Model Online Frame Num ScanNet++ Selected Subset Positive IoU
@@
MUSt3R >10,000>10{,}000 0.7168
@@
We compare 3D foundation models in Table 7. VGGT and $\pi^{3}$ operate offline on frame batches, unsuitable for online tracking. CUT3R supports online operation, but yields limited instance-level alignment as shown in Fig. 7, reflected by its Positive IoU of 0.2751 on the ScanNet++ Selected Subset. MUSt3R is fully online with substantially stronger object alignment across viewpoints. This alignment is essential: consistent 3D alignment enables stable 2D cross-view correspondence, allowing reliable mask propagation without explicit 3D merging.
@@ 5 Conclusion
We introduce 3AM, integrating 3D-aware features into SAM2 for viewpoint-robust video object segmentation. Through our Feature Merger and field-of-view aware sampling, 3AM achieves geometry-consistent tracking requiring only RGB at inference. On wide-baseline datasets, 3AM substantially outperforms state-of-the-art VOS methods, achieving gains of +15.9 and +30.4 points on the ScanNet++ Selected Subset.
Acknowledgements.
@@
Appendix A Overview
This supplementary material presents additional results to complement the main manuscript. First, we provide all the implementation details, including network architectures in App. B. Next, we elaborate on training details, such as the FoV sampling and datasets in App. C. Then, we describe how we perform class-agnostic instance segmentation in App. D. Finally, we provide more qualitative comparisons on both video object tracking and class-agnostic instance segmentation in App. E. In addition to this document, we provide additional video results to compare with state-of-the-art methods.
Appendix B Implementation Details
@@
B.1 Feature Merger
As illustrated in Fig. 4, our method uses multi-layer features from the MUSt3R decoder. We keep MUSt3R’s memory mechanism unchanged—i.e., we do not modify its view-coverage–based memory selection. We empirically select the layer indices $\text{encoder},4,7,11]$ to incorporate both semantic cues from early layers and geometry-aware signals from later layers, as deeper decoder layers in MUSt3R increasingly capture 3D structure and directly decode point clouds.
@@ B.1 Feature Merger
Appendix C Training Details
C.1 Field-of-View Sampling
@@ C.2 Settings
C.3 Datasets
@@
ScanNet++ is a high-quality indoor RGB–D reconstruction dataset that provides accurate camera poses, dense trajectories, and detailed geometric annotations. The public release contains over 1,000 reconstructed scenes, representing an expanded version of the initial smaller release[82]. Its combination of reliable geometry and viewpoint diversity makes it well suited for studying view-consistent object tracking under large camera motions. However, the 2D masks in ScanNet++ are generated by projecting 3D point-cloud instance labels into the image plane, which inevitably introduces reprojection noise from depth and pose inaccuracies. Despite this, ScanNet++ remains valuable because it provides realistic indoor scenes with rich geometric structure and large viewpoint variation. Together with the other two datasets, which supply clean, frame-aligned 2D masks, this forms a complementary training combination that balances realistic geometry with high-quality appearance supervision.
Appendix D Class-Agnostic Instance Segmentation
@@
Appendix E Qualitative Comparison
@@
We provide video object tracking visualization in Figs. 8, 9, 10, 11, 12, 13 and 14.
@@
We provide additional visual results on class-agnostic instance segmentation in Fig. 15.

### v5

@@ 4 Experiments
[…]
4.1.5 Replica.
On the Replica dataset, 3AM achieves the best performance across all metrics (Tab.˜4), reaching an IoU of 0.8119 and surpassing SAM2 (0.4424), SAM2Long (0.7691), and DAM4SAM (0.7744). It also attains the highest Tracking Recall (0.6381) and Accuracy (0.6793), demonstrating robust tracking under the large viewpoint variations characteristic of Replica.
These results demonstrate that 3AM consistently surpasses prior methods and is particularly effective in scenarios involving object reappearance and large viewpoint variation, despite not relying on the explicit memory-selection mechanisms used in previous approaches. Due to space limitations, more visualization is provided in the supplementary materials.
@@ 4.3 Ablation Study
4.3.1 Component Analysis.
We provide ablation studies in Tab.˜6 for Feature Source, Feature Merger Architecture, and Sampling Strategy. We evaluate variants using MUSt3R or SAM2 features alone under the same training protocol to demonstrate that both features are necessary. MUSt3R provides geometric grounding, while SAM2 captures appearance correspondence. Without MUSt3R, relying solely on appearance matching from SAM2 introduces ambiguity and leads to degraded performance. One thing to note is that SAM2 performs worse after finetuning. We attribute this degradation to the absence of grounding information from MUSt3R, which causes the memory attention to be mistrained due to insufficient reference cues for distinguishing different objects. Regarding the sampling strategy, naively applying random sampling results in worse performance than continuous sampling. When valid masks exist in both views, enforcing associations across large spatial discrepancies introduces ambiguity and leads to slower convergence.
Table 6: Component analysis on ScanNet++ selected subset.
Feature Source Feature Merger Sampling Strategy IoU Tracking Recall Accuracy
SAM2 & MUSt3R Addition & Convolution FoV Sampling 0.8771 0.6481 0.7149
SAM2 & MUSt3R Concatentation & Attention FoV Sampling 0.8831 0.6744 0.7462
SAM2 & MUSt3R Hierarchical Cross Attention (Final) Continuous Sampling 0.7925 0.6518 0.7513
SAM2 & MUSt3R Hierarchical Cross Attention (Final) Random Sampling 0.7363 0.5371 0.6311
MUSt3R - FoV Sampling 0.8461 0.5471 0.4129
SAM2 - FoV Sampling 0.6628 0.0038 0.0276
SAM2 & MUSt3R Hierarchical Cross Attention (Final) FoV Sampling 0.9061 0.7168 0.7737
4.3.2 Memory Selection.
Table 7: Comparison of different memory selection methods on ScanNet++ Whole Set.
Method IoU $\uparrow$ Tracking Recall $\uparrow$ Accuracy $\uparrow$
[…]
@@
Model Online Frame Num ScanNet++ Selected Subset Tracking Recall
@@
[…]
We further analyze the role of memory selection by combining 3AM with existing strategies proposed for SAM2. Table 8 reports results on the ScanNet++ Whole Set. 3AM, which uses the original SAM2 memory-selection mechanism, already achieves strong performance with an IoU of 0.8898. When we incorporate alternative memory-selection policies, such as those from DAM4SAM [77] or SAM2Long [18], we observe modest improvements: DAM4SAM-3AM slightly increases Accuracy to 0.7204, while SAM2Long-3AM achieves the highest IoU (0.9004) and Accuracy (0.7361). These results suggest that 3AM already delivers strong and stable performance without relying on specialized memory-selection policies. While alternative strategies such as those from SAM2Long or DAM4SAM offer small additional gains, their impact is relatively limited compared to the improvements brought by 3AM itself. We view the design of memory-selection mechanisms tailored specifically for 3AM as an interesting direction for future work.
4.3.3 3D Foundataion Models.
@@
We compare 3D foundation models in Table 8. VGGT and $\pi^{3}$ operate offline on frame batches, unsuitable for online tracking. CUT3R supports online operation, but yields limited instance-level alignment as shown in Fig.˜7, reflected by its Tracking Recall of 0.2751 on the ScanNet++ Selected Subset. MUSt3R is fully online with substantially stronger object alignment across viewpoints. This alignment is essential: consistent 3D alignment enables stable 2D cross-view correspondence, allowing reliable mask propagation without explicit 3D merging.
@@ 5 Conclusion
We introduce 3AM, integrating 3D-aware features into SAM2 for viewpoint-robust video object segmentation. Through our Feature Merger and field-of-view aware sampling, 3AM achieves geometry-consistent tracking requiring only RGB at inference. On wide-baseline datasets, 3AM substantially outperforms state-of-the-art VOS methods, achieving gains of +15.9 and +30.4 points on the ScanNet++ Selected Subset, while automatically switching back to SAM2’s capability to handle dynamic video scenarios.
Acknowledgements
@@
Appendix 0.A Overview
This supplementary material presents additional results to complement the main manuscript. First, we provide all the implementation details, including network architectures in Appendix˜0.B. Next, we elaborate on training details, such as the FoV sampling and datasets in Appendix˜0.C. Then, we describe how we perform class-agnostic instance segmentation in Appendix˜0.D. Finally, we provide more qualitative comparisons on both video object tracking and class-agnostic instance segmentation in Appendix˜0.E. In addition to this document, we provide additional video results to compare with state-of-the-art methods.
Appendix 0.B Implementation Details
@@
0.B.1 Feature Merger
As illustrated in Fig.˜4, our method uses multi-layer features from the MUSt3R decoder. We keep MUSt3R’s memory mechanism unchanged—i.e., we do not modify its view-coverage–based memory selection. We empirically select the layer indices $\text{encoder},4,7,11]$ to incorporate both semantic cues from early layers and geometry-aware signals from later layers, as deeper decoder layers in MUSt3R increasingly capture 3D structure and directly decode point clouds.
@@ 0.B.1 Feature Merger
Appendix 0.C Training Details
0.C.1 Field-of-View Sampling
@@ 0.C.2 Settings
Prompts
Due to the degraded 2D mask and the large viewpoint variations introduced by the sampling strategy during training, we restrict the input modality to masks only for ScanNet++ and ASE. For MOSE, we enable all prompt types, including point, box, and mask inputs.
0.C.3 Datasets
@@
ScanNet++ is a high-quality indoor RGB–D reconstruction dataset that provides accurate camera poses, dense trajectories, and detailed geometric annotations. The public release contains over 1,000 reconstructed scenes, representing an expanded version of the initial smaller release[109]. Its combination of reliable geometry and viewpoint diversity makes it well suited for studying view-consistent object tracking under large camera motions. However, the 2D masks in ScanNet++ are generated by projecting 3D point-cloud instance labels into the image plane, which inevitably introduces reprojection noise from depth and pose inaccuracies. Besides, the pose annotations in ScanNet++ often omit long stretches of consecutive frames, as the SfM pipeline discards uncertain frames. Under such conditions, our FoV-based sampling becomes even more important. Despite this, ScanNet++ remains valuable because it provides realistic indoor scenes with rich geometric structure and large viewpoint variation. Together with the other two datasets, which supply clean, frame-aligned 2D masks, this forms a complementary training combination that balances realistic geometry with high-quality appearance supervision.
Appendix 0.D Class-Agnostic Instance Segmentation
@@
Appendix 0.E Qualitative Comparison
@@
We provide video object tracking visualization in Figs.˜8, 9, 10, 11, 12, 13 and 14.
@@
We provide additional visual results on class-agnostic instance segmentation in Fig.˜15.